1.if 1/5^3y=0.008,then find the value of 0.25^y

Grade:12

2 Answers

Aman Bansal
13 years ago

DEAR PANKAJ,

if 5-3y= (2*10-1)3

=> 5-y = 2*10-1

=> 5^y = 1/2*10

=>5^y= (5)^1

=>y=1

0.25^y=0.25^1=0.25

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

AMAN BANSAL

Fawz Naim
13 years ago

Here it is given that   1/5^3y=0.008

=>1/5^3y=8/1000

=>5^3y=1000/8    (taking reciprocal on both sides)

5^3y=(10/2)^3

5^3y=5^3

as base on both the sides are equal then the powers will also be equal

=>3y=3

=>y=1   now 0.25^y=0.25^1=0.25